OpenStudy (anonymous):

im so lost :/ hellppppp pleasee! Write a two-column proof. Given: 7y = 8x – 14; y = 6 Prove: x = 7

OpenStudy (anonymous):

Ok let us check....... Substitute y = 6 you get \[7 \times 6 =8x -14\] \[ \rightarrow 42+14 = 8x\] \[ \rightarrow 56 = 8x\] \[ \rightarrow x = \frac{ 56 }{ 8 } = 7\]

OpenStudy (anonymous):

thus your asnwer. is proved.... :)

OpenStudy (anonymous):

wait im lost.. after the first part lol.. so how did you get 42? on the second part? im honestly confused. and it has to be atwo column.. meaning one side has a statement and the other side has.. given.. or something like that. :/

OpenStudy (anonymous):

here in your question Y = 6 .... that means that to find a solution both equation has to satisfy.. means you have to find a value of 'x' when 'y' is equal to 6... so substitute the value of Y (=6) in 7y = 8x – 14...

OpenStudy (anonymous):

okay i get that part.. now the other part of the question is for each step.. your suppose to state what method you use.. do you know what i mean?

OpenStudy (anonymous):

im sorry this is difficult for me.. im a little slow :p

OpenStudy (anonymous):

lol ...sorry .lol ok \[7 \times 6 = 42\]

OpenStudy (anonymous):

your killin me lol. okay ill explain what im trying to say.. in a two column proof. it looks a little like this.. Given: <2 is congruent to <4, M<2=110 Proof: m<3 = 70 m∠2 = m∠4 Definition of congruent angles m∠4 = 110º b. ________ ∠3 and ∠4 are a linear pair Definition of a linear pair (shown in diagram) ∠3 and ∠4 are supplementary Linear Pair postulate m∠3 + m∠4 = 180º c. _______ m∠3 = 70º d. ________ (3 points)

OpenStudy (anonymous):

idek if im making any sense right now im trying to help you understand wht i mean :/ lol

OpenStudy (anonymous):

Thank you But can I ask is there any defenition like that... on ly thing is algbic manipulation... th concept be hind this question is explaind in my 3rd answer. I am still confused How the two colum proof can be given in your question since it is only a algebric manipulation.. Do you want me to explain each algbric manipulation on the othr side......??????..I wounder for what step you want explation

OpenStudy (anonymous):

yeah thats exactly what i mean, for each step, beside it, you write what step you used, like subtraction property of equality etc etc. thank you soo much for helping me on this!

OpenStudy (anonymous):

and like you said for answer 3 i think that one would be substitution property of equality right?

OpenStudy (anonymous):

yea... It is ..... but we used that property to find the solution of equation... means the point which satisfies both equation

OpenStudy (anonymous):

yes, but what about all the other steps?

OpenStudy (anonymous):

im soo giving you a medal !!!

OpenStudy (anonymous):

..in other steps you add both sides + 14... that is adding or subtracting or multiplying or dividing both sides of equality by a number won't affct the value

OpenStudy (anonymous):

can you write each one for each step out? please

OpenStudy (anonymous):

7×6=8x−14 →42+14=8x-14+14 (using the above proprty ... let us say invariabilty of equality on addition on both sides)) →56=8x (addition of two numbrs) \[\frac{ 56 }{8 } = \frac{x8 }{ 8 }\] (usig the above propety... let us say invariabilty of equality on division on both sides) \[x =\frac{ 56 }{ 8 }\] \[ x = 7\] by division ..

OpenStudy (anonymous):

thank you so much! could yuo help me with one more? its alot easier lol

OpenStudy (anonymous):

oh, god... It was a bitter task.. but it helped me to know two column explanation :)

OpenStudy (anonymous):

it was hard. this online schooling is difficult at times

OpenStudy (anonymous):

What is M... and x reprsnt.... Angles... I studied a long time back.. can you xplain

OpenStudy (anonymous):

okay we'll do this one together, m equals the measurment of that particular angle. its easier to draw it out i just get stuck on the equation :/ lol

OpenStudy (anonymous):

and im not sure how to set the equation up...

OpenStudy (anonymous):

OK... now I got some Idea.... since BD bisects angle ABC it will b som thing like this since angle is bisctd .. each angle(ABD,DBC) will be equal which is Y..(bisection Property) But from your question... 2Y = 8x (since measurement of ABC = 8x) and also from your question measurement of ABD = 2x+22 = Y; But we know 2Y = 8x .. there for Y = 4x...(division) therefore 2x+22 =4x ( substitution of equality-- equal values can be replavd with out chaging the fuunction) 22+2x-2x=4x-2x (let us say invariabilty of equality on subtraction on both sides) 22 = 2x (subtraction) x = 11 (division) But we know Y = 4x, by substitution of quality, we get \[y= 4 \times 11\] = 44 which is the value of DBC since Y =since measurement of DBC too

OpenStudy (anonymous):

And valu of measure ment of angle ABC = 2Y,by substitution of equality and multiplication we get.... It is \[2 \times 44 = 88\]